# Inferring Capabilities from Task Performance with Bayesian Triangulation

### v1

@@ D.3 Object Permanence Measurement Layout
OPAbility Capability $[0,48.4]$ $Uniform(0,48.4)$
@@
rampAbility Capability $[0,1]$ $Beta(1,1)$
lavaAbility Capability $[0,1]$ $Beta(1,1)$
platformAbility Capability $[0,1]$ $Beta(1,1)$
@@
rightleftBias Bias [,][$\infty$,$\infty$][ - ∞ , ∞ ] $\mathcal{N}(0,1)$
@@
rightleftEffect INON [,][$\infty$,$\infty$][ - ∞ , ∞ ] ${\sf rightLeftBias}\times{\sf rightLeftPosition}$
@@
Table 5 gives the complete details, broken down by node type. As in the AAIO case, the function $\sigma$ is the standard logistic function while $\omega$ is a parameterised weighting function, where the value $\nu$ represents a noise prior that represents a constant model set to 1 -- average performance. The function $\mu$ is a margin, defined as
@@
D.4 Synthetic Agents and Comparison
The synthetic agents Team A designed can be found in Table  6 and Table  7.
@@ D.4 Synthetic Agents and Comparison
Agent 1
Reference
Passes every instance.
Agent 2
Reference
Passes each instance with a probability of 0.5.
Agent 3
Reference
Passes each instance with a probability of 0.1.
Agent 4
Reference
Approximates a random walker. Passes a small proportion of instances. It is more likely to pass ’safe’ instances that don’t have lots of dangerous holes to fall down or lava to run into.
Agent 5
Reference
Passes each instance with a probability of 0.9. This is a more human-like failure rate than agent 1.
Agent 6
Achilles Heel
Has OP, passing an instance with a probability of 0.95 unless there is lava present, in which case it passes with a probability of 0.1.
Agent 7
Achilles Heel
Has OP, passing an instance with a probability of 0.95 unless there is lava or an occluder with a ‘redness‘ over 200 present, in which case it passes with a probability of 0.1.
Agent 8
Achilles Heel
Has OP, passing an instance with a probability of 0.95 unless there is an occluder with a ‘greenness‘ over 200 present, in which case it passes with a probability of 0.1.
Agent 9
Achilles Heel
Has OP, passing an instance with a probability of 0.95 unless there is a ramp present, in which case it passes with a probability of 0.1.
Agent 10
Achilles Heel
Has OP but lacks fine control of actions. If there is lava present, it only passes those instances with a probability of 0.5. It often falls into the wrong holes on its way to the further holes, so as distance to goal increases, the probability of passing Grid instances increases from 0.6 in increments of 0.1. This is only for the OP tasks. For the non-OP instances and the remaining instances, it solves an instance with a probability of 0.8.
Agent 11
Achilles Heel
Has OP but poor visual acuity. If goals are 0.5 or smaller, they only pass the instance with a probability of 0.1. For goals of size 1, the probability is 0.3, it is 0.5 for size 1.5, 0.7 for size 2, 0.8 for size 2.5, and 0.9 for a goal size 3 or larger.
Agent 12
Achilles Heel
Has OP but is biased to the right. Probability of passing a task is 0.95 when the goal is to the right, 0.3 when it is to the left, and 0.6 when it is centrally located.
Agent 13
Achilles Heel
Has OP but is biased to the left. Probability of passing a task is 0.95 when the goal is to the left, 0.3 when it is to the right, and 0.6 when it is centrally located.
Agent 14
Achilles Heel
Has OP, but is easily confused when there are several positions that are similar looking (either identical or mirror images of each other).
Agent 15
Fraudster
Does not have OP, but goes to where they last saw the reward. Deterministically passes all instances where this policy works (all tasks except some of CV Chick tasks), and fails the rest.
Agent 16
Fraudster
Does not have OP, but goes to where they last saw the reward. Passes all instances where this policy works with probability of 0.95 (all tasks except some of CV Chick tasks), and passes the rest with probability 0.25.
Table 6: The agents Team A designed and synthesised performances for.
@@
Agent 17
Achilles Heel
@@
Has OP, but has low memory, so struggles when the goal is occluded for longer.
Agent 18
Achilles Heel
@@
Has OP, but is easily confused when there are more positions that the goal could be occluded in. Passes instances with more than 10 other positions available with probability of 0.3. Between 5 and 10, probability of 0.6. Between 3 and 5, probability of 0.75. Less than 3, passes with probability of 0.95.
Agent 19
Context-Specific
@@
Non-robust OP. It can solve the navigation tasks and all the CV tasks, but fails the other instances.
Agent 20
Context-Specific
@@
Non-robust OP. It solves the navigation tasks and all the CV tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 21
Context-Specific
@@
Non-robust OP. It can solve the navigation tasks and all the Grid tasks, but fails the other instances.
Agent 22
Context-Specific
@@
Non-robust OP. It can solve the navigation tasks and all the Grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 23
Context-Specific
@@
Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks, but fails the other instances.
Agent 24
Context-Specific
@@
Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 25
Context-Specific
@@
Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks, but fails the other instances.
Agent 26
Cognitive Pathology
@@
Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 27
Cognitive Pathology
@@
Non-robust OP. It can solve the navigation tasks and all the 3 Cup and grid tasks, but fails the other instances.
Agent 28
Cognitive Pathology
@@
Non-robust OP. It can solve the navigation tasks and all the 3 cup and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Agent 29
Cognitive Pathology
@@
Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks, but fails the other instances.
Agent 30
Cognitive Pathology
@@
Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Table 7: The agents Team A designed and synthesised performances for (continued).
@@
Figs. 13 and 14 show the models inferred values for each element of the cognitive profile. Fig 13 gives the mean of the inferred distribution and Fig .14 the standard deviation. These values were used by Team B to infer the qualitative type of agent being evaluated. Multiple aspects of the values inferred by the model need to be considered to make a proper assessment of each agent—particularly since Team B did not know a priori the exact types of agents that would exist. For example, Agent 2 has relatively low scores for most abilities, a high standard deviation for most abilities, and a value of $0.5$ for the mean and s.d of overall success. This led Team B to believe that this agent was succeeding the tasks at random with 50% probability. Equally, when looking at Agents 13 and 14, the right-left bias was high (with low standard deviation) leading Team B to believe that these agents were massively favouring movement in particular directions.
@@
Team B were able to partially recover the cognitive profiles of the context-specific agents. The relatively lower accuracy here was because task type was not included in the measurement layout. Therefore, many of the biases recovered by Team B reflect the particular construction of the paradigms used by Team A (see legend of Fig.  15).
Team B did not accurately recover the cognitive profiles of the ‘fraudster’ agents. Both Agents 15 and 16 follow the policy of going to where the goal was last seen. This results in generally good performance in all tasks except the Chiandetti & Vallortigara tasks, where this policy would result in falling into lava. Therefore, the performance distributions look similar to a high performing object permanence agent. To successfully discriminate these sophisticated agents, we would need to weight the importance of certain sub-groups of instances, such that, for example, Chiandetti & Vallortigara tasks are of most inferential importance in the case where there is good performance on the remaining paradigms. This requires targeted inference and diligent use and application of measurement layouts.
Overall, however, Team B was able to extract a lot of information about a the majority of the synthetic agents—despite an imperfectly designed measurement layout. Also noteworthy is that when the measurement provides correct inferences, they tend to be consistent across that particular agent. That is, if the measurement layout identifies a particular weakness in an agent, it won’t also pick a second, nonexistent weakness also.
@@
Appendix E Prediction Results
@@
This section gives the full results for prediction on unseen instances. The synthetic data set was partitioned into an 80/20 train-test split. The measurement layout models (one for each agent) were trained on the same training partition. For each instance from the test-set we apply forward inference to yield a success probability.
@@
We evaluate these probabilities using a Brier score (Brier 1950), also making use of the decomposition into Calibration and Refinement (Murphy 1973). These scores are given in Table 8, with the better performing approach for prediction in bold for each agent. As expected, the aggregate predictions have no refinement at all, and the values are close to 0.25 for balanced cases (the lower numbers only appear when the proportion of right/wrong performance is more imbalanced). The results from the measurement layout are much better here, as this method can give refined results per instance. When we look at calibration, however, we see aggregate predictions are almost perfect, since train and test come from the same distribution. But the calibration from the measurement layout is also quite good, and overall there are many more wins for the measurement layouts approach. Moreover, it is important to emphasise that unlike this table, train and test do not have to come from the same distribution. For instance, if we removed the easy instances for many of the tasks, we would get similar results for the measurement layout prediction approach (calibration and refinement should not change dramatically), but the results for the aggregate prediction approach would be much worse, since the calibration would be completely lost.
Looking at the particular agent properties that cause the measurement layouts to be more or less capable at prediction can provide some insight into the strengths and weaknesses of our approach. Agent 1 was designated as the perfect agent that will always pass every instance. It is no surprise then that predictions using the aggregate success rate were very successful, achieving the best possible Brier score. Our model’s predictions were still highly accurate (Brier score of 0.022), but the model lacked the extreme certainty of the aggregate measure—likely because using Bayes’ rule, beginning with an uninformed prior it’s impossible to update the posterior to exclude one of the outcomes.
@@
Agent 5 is another case where the aggregate is better at prediction than our measurement layouts. Agent 5 passed all instances with probability 0.9. We believe our model struggled to perform as well as the aggregate here due to there not being sufficiently many instances to move the initial prior (uniform distribution) to the more extreme distribution predicting success at a rate of 0.9. Our model was able to cope with other fixed pass rates (such as with Agent 2, 3, and 4), so this may just be a matter of degree. It’s worth noting that for any fixed pass rate Agent, it’s actually impossible to predict more accurately than the aggregate in expectation. Hence For these types of agents the Aggregate Brier score is minimised, and for many fixed pass rate agents, our measurement layouts matched the aggregate’s Brier score.
@@
The only other agents that our measurement layouts were worse at predicting than the aggregate were agents 15 and 16. These two agents were designed as ‘Fraudster’ agents that mimic object permanence by going to the last position where they saw the reward. This is sufficient to solve many object permanence tasks except for some of the CV chick tasks. When this ‘Fraudster’ policy would be successful, these agents succeed with a probability of 0.95, if the policy would not be applicable, the agents succeeds with probability 00 (for agent 15) or $0.25$ (for agent 16). We believe it is the high overall success rate that causes our model to be less predictive here, similar to Agent 5 (and to a lesser extent Agent 7 and 8), again because it takes more task instances to move an uninformative policy to extremes such as predicting success with a probability greater than $0.9$.

### v2

@@ D.2 Experiment 2: An Object Permanence Task
OPAbility Capability $[0,48.4]$ $Uniform(0,\sf memoryAbility\times maxPositions)$
@@
rampAbility Capability $[0,1]$ $Uniform(0,1)$
lavaAbility Capability $[0,1]$ $Uniform(0,1)$
platformAbility Capability $[0,1]$ $Uniform(0,1)$
@@
rightleftBias Bias $\infty$$\infty$] $\mathcal{N}(0,1)$
@@
rightleftEffect INON $\infty$$\infty$] ${\sf rightLeftBias}\times{\sf rightLeftPosition}$
@@
The function $\sigma$ is the standard logistic function while $\omega$ is a parameterised weighting function, where the value $\nu$ represents a noise prior that represents a constant model set to 1 - average performance. The function $\mu$ is a margin, defined as
@@
D.3 Experiment 3: Extending Measurement Layouts To Real Data
To accommodate the data from Voudouris et al. (2024), we updated the measurement layout, which is presented in Figure 16.
Figure 16: The updated measurement layout to accommodate the meta-features and observed variables collected in Voudouris et al. (2024).
There are three central capabilities: OP, navigation, and visual acuity. There are four further capabilities that are pertinent to navigation: the capability to avoid death zones, and the capabilities to navigate to goals placed to the left, directly ahead, and to the right of the agent’s starting position. A full description is given in Table 4 including all the value ranges, distributional priors and linking functions.
Name Type Range/Values Prior/Linking Function
Number of Positions Meta-feature $\{0,12\}$
Goal Occluded Meta-feature $\{0,1\}$
Goal Distance Meta-feature $[9,118]$
Number of Turns To Goal Meta-feature $[0,13]$
Choice Distance Meta-feature $[1,53.5]$
Number of Turns To Choice Meta-feature $[0,6]$
Goal Size Meta-feature $[0.5,5]$
Death Zone Presence Meta-feature $\{0,1\}$
Goal Ahead Meta-feature $\{0,1\}$
Goal Left Meta-feature $\{0,1\}$
Goal Right Meta-feature $\{0,1\}$
OPAbility Capability $0,\infty]$ $\text{HalfNormal}(503.1)$
Flat Navigation Ability Capability $[0,4.4]$ $\text{HalfNormal}(163.80)$
Visual Acuity Capability $0,\infty]$ $\text{HalfNormal}(47.80)$
Lava Ability Capability $0,\infty]$ $\text{HalfNormal}(1)$
Ahead Ability Capability $0,\infty]$ $\text{HalfNormal}(1)$
Right Ability Capability $0,\infty]$ $\text{HalfNormal}(1)$
Left Ability Capability $0,\infty]$ $\text{HalfNormal}(1)$
OPSuccessPerformance INON $[0,1]$ $\sigma({\sf OPAbility},{\sf GoalOccluded*(GoalDistance-ChoiceDistance)*NumberOfPositions})$
OPChoicePerformance INON $[0,1]$ $\sigma({\sf OPAbility},{\sf GoalOccluded*ChoiceDistance*NumberOfPositions})$
FlatNavGoalPerformance INON $[0,1]$ $\sigma({\sf FlatNavAbility},{\sf GoalDistance*NumberTurnsGoal})$
FlatNavChoicePerformance INON $[0,1]$ $\sigma({\sf FlatNavAbility},{\sf ChoiceDistance*NumberTurnsChoice})$
visualAcuityPerformance INON $[0,1]$ $\sigma({\sf visualAbility}-\text{log}({\sf goalDistance}/{\sf goalSize}))$
lavaPerformance INON $[0,1]$ $\mu({\sf lavaAbility},{\sf lavaPresence})$
AheadPerformance INON $[0,1]$ $\mu({\sf AheadAbility},{\sf AheadPresence})$
LeftPerformance INON $[0,1]$ $\mu({\sf LeftAbility},{\sf LeftPresence})$
RightPerformance INON $[0,1]$ $\mu({\sf RightAbility},{\sf RightPresence})$
SuccessPerformance Observed $[0,1]$ $p\sim Bernoulli({\sf FlatNavGoalPerformance}\times{\sf OPSuccessPerformance}\times{\sf visualAcuityPerformance}\times{\sf lavaPerformance})$
ChoicePerformance Observed $[0,1]$ $p\sim Bernoulli({\sf FlatNavChoicePerformance}\times{\sf OPChoicePerformance}\times{\sf visualAcuityPerformance}\times{\sf AheadPerformance}\times{\sf LeftPerformance}\times{\sf RightPerformance})$
Appendix E Evaluating the Presence of Object Permanence
Imagine a simple task where a reward is observed to roll leftwards behind an occluder, and go fully out of sight. An agent with OP ought to pass this task, since it: (a) wants to obtain the reward, (b) understands that the reward continues to exist despite not being able to see it, and (c) is able to infer the location of the occluded reward based on its previously observed trajectory. However, an agent with OP might fail this task, if for instance it is bad at navigating, and takes a very circuitous route to its goal and runs out of time before it can reach the reward. Conversely, an agent lacking OP might pass this task, simply because it has been rewarded previously for going forwards some number of steps and then turning left. This is the problem of the underdetermination of theory by behaviour (Stanford, 2006), in which an observed behaviour can be explained in multiple ways. The use of well-informed measurement layouts based on internally valid test batteries allows us to more robustly eliminate incorrect alternative explanations and triangulate on the underlying causal explanations of AI system behaviour.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
E.1 Synthetic Agents and Comparison
The synthetic agents Team A designed can be found in Table  5, Table 6, and Table 7.
@@ E.1 Synthetic Agents and Comparison
Agent 1 Reference Passes every instance.
Agent 2 Reference Passes each instance with a probability of 0.5.
Agent 3 Reference Passes each instance with a probability of 0.1.
Agent 4 Reference Approximates a random walker. Passes a small proportion of instances. It is more likely to pass ’safe’ instances that don’t have lots of dangerous holes to fall down or lava to run into.
Agent 5 Reference Passes each instance with a probability of 0.9. This is a more human-like failure rate than agent 1.
Agent 6 Achilles Heel Has OP, passing an instance with a probability of 0.95 unless there is lava present, in which case it passes with a probability of 0.1.
Agent 7 Achilles Heel Has OP, passing an instance with a probability of 0.95 unless there is lava or an occluder with a ‘redness‘ over 200 present, in which case it passes with a probability of 0.1.
Agent 8 Achilles Heel Has OP, passing an instance with a probability of 0.95 unless there is an occluder with a ‘greenness‘ over 200 present, in which case it passes with a probability of 0.1.
Agent 9 Achilles Heel Has OP, passing an instance with a probability of 0.95 unless there is a ramp present, in which case it passes with a probability of 0.1.
Agent 10 Achilles Heel Has OP but lacks fine control of actions. If there is lava present, it only passes those instances with a probability of 0.5. It often falls into the wrong holes on its way to the further holes, so as distance to goal increases, the probability of passing Grid instances increases from 0.6 in increments of 0.1. This is only for the OP tasks. For the non-OP instances and the remaining instances, it solves an instance with a probability of 0.8.
Table 5: The agents Team A designed and synthesised performances for. Agents 1-10.
@@
Agent 11 Achilles Heel Has OP but poor visual acuity. If goals are 0.5 or smaller, they only pass the instance with a probability of 0.1. For goals of size 1, the probability is 0.3, it is 0.5 for size 1.5, 0.7 for size 2, 0.8 for size 2.5, and 0.9 for a goal size 3 or larger.
Agent 12 Achilles Heel Has OP but is biased to the right. Probability of passing a task is 0.95 when the goal is to the right, 0.3 when it is to the left, and 0.6 when it is centrally located.
Agent 13 Achilles Heel Has OP but is biased to the left. Probability of passing a task is 0.95 when the goal is to the left, 0.3 when it is to the right, and 0.6 when it is centrally located.
Agent 14 Achilles Heel Has OP, but is easily confused when there are several positions that are similar looking (either identical or mirror images of each other).
Agent 15 Fraudster Does not have OP, but goes to where they last saw the reward. Deterministically passes all instances where this policy works (all tasks except some of CV Chick tasks), and fails the rest.
Agent 16 Fraudster Does not have OP, but goes to where they last saw the reward. Passes all instances where this policy works with probability of 0.95 (all tasks except some of CV Chick tasks), and passes the rest with probability 0.25.
@@
Agent 17 Achilles Heel Has OP, but has low memory, so struggles when the goal is occluded for longer.
@@
Agent 18 Achilles Heel Has OP, but is easily confused when there are more positions that the goal could be occluded in. Passes instances with more than 10 other positions available with probability of 0.3. Between 5 and 10, probability of 0.6. Between 3 and 5, probability of 0.75. Less than 3, passes with probability of 0.95.
@@
Agent 19 Context-Specific Non-robust OP. It can solve the navigation tasks and all the CV tasks, but fails the other instances.
@@
Agent 20 Context-Specific Non-robust OP. It solves the navigation tasks and all the CV tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Table 6: The agents Team A designed and synthesised performances for. Agents 11-20.
Name Agent Class Agent Description
@@
Agent 21 Context-Specific Non-robust OP. It can solve the navigation tasks and all the Grid tasks, but fails the other instances.
@@
Agent 22 Context-Specific Non-robust OP. It can solve the navigation tasks and all the Grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
@@
Agent 23 Context-Specific Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks, but fails the other instances.
@@
Agent 24 Context-Specific Non-robust OP. It can solve the navigation tasks and all the 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
@@
Agent 25 Context-Specific Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks, but fails the other instances.
@@
Agent 26 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the CV and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
@@
Agent 27 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the 3 Cup and grid tasks, but fails the other instances.
@@
Agent 28 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the 3 cup and grid tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
@@
Agent 29 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks, but fails the other instances.
@@
Agent 30 Cognitive Pathology Non-robust OP. It can solve the navigation tasks and all the CV and 3 cup tasks with a probability of 0.75, fails the other paradigms with probability of 0.1.
Table 7: The agents Team A designed and synthesised performances for. Agents 21-30.
E.2 Object Permanence Qualitative Results
@@
Figs. 19 and 20 show the models inferred values for each element of the cognitive profile. Fig 19 gives the mean of the inferred distribution and Fig .20 the standard deviation. These values were used by Team B to infer the qualitative type of agent being evaluated. Multiple aspects of the values inferred by the model need to be considered to make a proper assessment of each agent—particularly since Team B did not know a priori the exact types of agents that would exist. For example, Agent 2 has relatively low scores for most abilities, a high standard deviation for most abilities, and a value of $0.5$ for the mean and s.d of overall success. This led Team B to believe that this agent was succeeding the tasks at random with 50% probability. Equally, when looking at Agents 13 and 14, the right-left bias was high (with low standard deviation) leading Team B to believe that these agents were massively favouring movement in particular directions. We also present Forest Plots for agents, showing their 95% HDI across the 8 core capabilities in Figures 17 and 18.
(a) Object permanence
(b) Flat navigation
(c) Visual acuity
(d) Lava
Figure 17: 95% HDIs for all agents across Object Permanence, Flat Navigation, Visual Acuity, and Lava.
(a) Platform
(b) Ramp
(c) Right Left Bias
(d) Memory
Figure 18: 95% HDIs for all agents across Platforms, Ramps,
@@
Team B were able to partially recover the cognitive profiles of the context-specific agents. The relatively lower accuracy here was because task type was not included in the measurement layout. Therefore, many of the biases recovered by Team B reflect the particular construction of the paradigms used by Team A (see legend of Figure 21).
Team B did not accurately recover the cognitive profiles of the ‘fraudster’ agents. Both Agents 15 and 16 follow the policy of going to where the goal was last seen. This results in generally good performance in all tasks except the Chiandetti & Vallortigara tasks, where this policy would result in falling into lava. When this ‘Fraudster’ policy would be successful, these agents succeed with a probability of 0.95, if the policy would not be applicable, the agents succeeds with probability 0 (for agent 15) or $0.25$ (for agent 16). Therefore, the performance distributions look similar to a high performing object permanence agent and our approach attributed higher values of OP than were warranted. To successfully discriminate these sophisticated agents, we would need to weight the importance of certain sub-groups of instances, such that, for example, Chiandetti & Vallortigara tasks are of most inferential importance in the case where there is good performance on the remaining paradigms. This requires good alignment between those designing the task and those creating the measurement layouts to make sure that controls established in the task design are reflected in the analysis.
Overall, however, Team B was able to extract a lot of information about a the majority of the synthetic agents—despite an imperfectly designed measurement layout. Also noteworthy is that when the measurement provides correct inferences, they tend to be consistent across that particular agent. That is, if the measurement layout identifies a particular weakness in an agent, it won’t also pick a second, nonexistent weakness.
Capability OP FlatNav Visual Lava Platform Ramp Memory
Overall RMSE 0.28 0.18 0.16 0.20 0.28 0.38 0.15
Included RMSE 0.13 0.11 0.24 0.17 0.13 0.27 0.16
@@
E.3 Object Permanence Predictive Results
@@
The synthetic dataset was partitioned into an 80/20 train-test split. The measurement layout models (one for each agent) were trained on the same training partition. For each instance from the test-set we apply forward inference to yield a success probability.
We evaluate these probabilities using a Brier score (Brier, 1950), also making use of the decomposition into Calibration and Refinement (Murphy, 1973). Figure 5(a) plots Brier score against average success rate. The measurement layout model is consistently a better predictor of success than relying on the aggregate measure. We see better prediction performance for agents with less certain success rates. These are the agents for which we are more interested in improving prediction as there is more capacity to reduce uncertainty.
@@
These scores are also given fully in Table 9, with the better performing approach for prediction in bold for each agent. As expected, the aggregate predictions have no refinement at all, and the values are close to 0.25 for balanced cases (the lower numbers only appear when the proportion of right/wrong performance is more imbalanced). The results from the measurement layout are much better here, as this method can give refined results per instance. When we look at calibration, however, we see aggregate predictions are almost perfect, since train and test come from the same distribution. But the calibration from the measurement layout is quite high, and overall there are many more cases where the measurement layout approach is superior than the aggregate. Moreover, it is important to emphasise that unlike this table, train and test do not have to come from the same distribution. For instance, if we removed the easy instances for many of the tasks, we would get similar results for the measurement layout prediction approach (calibration and refinement should not change dramatically), but the results for the aggregate prediction approach would be much worse, since the calibration would be completely lost. However, doing this
Looking at the particular agent properties that result in the measurement layouts to be more or less capable at prediction can provide some insight into the strengths and weaknesses of our approach.
@@
Agents 1 and 5 were designed to pass all instances with probability 1 and 0.9 respectively. It’s no surprise that predictions using the aggregate success rate were very successful. Our model’s predictions were still highly accurate (Brier scores of 0.022 and 0.099 respectively), but the model lacked the extreme certainties of the aggregate measures. For Agent 1, this would be impossible—using Bayes’ rule and an uninformed prior it’s not possible to update the posterior to completely exclude one of the outcomes. For agent 5 we believe that our model didn’t achieve the same level of certainty of success due to there not being sufficiently many instances to move the initial prior (uniform distribution) to the more extreme distribution predicting success at a rate of 0.9. Our model was able to cope with other fixed pass rates (such as with agents 2, 3, and 4), so this may just be a matter of degree. It’s worth noting that for any fixed pass rate agent, it’s actually impossible to predict more accurately than the aggregate in expectation. Hence for these types of agents the Aggregate Brier score is minimised, and for many fixed pass rate agents, our measurement layouts matched the aggregate’s Brier score.
@@
The only other agents that our measurement layouts were worse at predicting than the aggregate were agents 15 and 16. These two agents were designed as ‘Fraudster’ agents that mimic object permanence by going to the last position where they saw the reward. We discussed these agents and their strategy in the Section 4 about Qualitative Validity. Since these agents had very high overall success rates, we believe that this causes our model to be less predictive, similar to Agent 5 (and to a lesser extent Agent 7 and 8), because it takes more task instances to move an uninformative policy to extremes such as predicting success with a probability greater than $0.9$.